# Topological Spatial Graph Coarsening

### v1

@@ 3.2 Persistence Diagram for Spatial Graphs
We want to adapt the above construction of persistence diagrams of point clouds data to spatial graph data. To do so, we follow the same general idea: build a nested family of simplicial complexes on top of graphs and monitor how the topological components evolve through this filtration.
Previous approaches for computing filtrations over graphs considered a family of nested subgraphs of $G$, $(G_{r})_{r\geq 0}$ (by filtering the edges based on weights or other graph characteristics like degree) and compute simplicial complexes over each subgraph $C(G_{r})$ where $C$ can be for example the clique complex (also called flag complex) or the neighborhood complex [lovasz1978kneser, aharoni2005eigenvalues, horak2009persistent]. These approaches differ from the point cloud pipeline on different aspects. First, the scale is given by a substructure of the initial object (here the graph) while for point clouds we consider the whole dataset and compute simplicial complexes at different scales. Secondly, some simplices are never included in the filtration (e.g., a 1-simplex that correspond to an edge that is not in $G$ will never be include in the clique complex).
Hence, we propose a new way of computing a filtration of simplicial complexes over a graph that is closer to the construction for point clouds. The essence of our approach lies in the fact that the VR filtration is entirely determined by the distances between all pairs of points in the data. Thus, we propose to replace the euclidean distance by the shortest path distance over the graph, denoted by $d_{G}$ and defined for all pairs of vertices $u$ and $v$ by
@@
where we recall that the $\ell_{u,v}$ of an edge $(u,v)$ is given by the euclidean distance $d(x_{u},x_{v})$. Therefore, we define a simplicial complex at scale $r>0$ over the graph $G$ as $\mathcal{C}_{r}(G)=VR_{r}(V,d_{G})$.
@@
If the graph is the complete graph, the shortest path distance coincides with the euclidean distance and we recover the classical $V_{r}$ filtration.
The proposed filtration $\mathcal{C}(G)=(\mathcal{C}_{r}(G))_{r\geq 0}$ has the drawback that cycles formed by a triplet of edges (empty triangle) do not persist in the filtration because the 2-simplex (filled triangle) appears in the filtration exactly when the longest edge is included, making the cycle disappear at the same scale at which it appears. Thus, this type of cycles are not represented in the induced persistence diagram. As we would like to represent in the persistence diagram any cycles - no matter the number of edges - we propose to modify the construction of the simplicial complexes $\mathcal{C}_{r}(G)$ by enlarging the scale at which 2-simplices appear.
@@ 4 Topological Spatial Graph Coarsening
Given a sequence of coarsened graphs obtained by exploring the set $\Theta$, we define a scoring function to select the optimal coarsened graph. In graph coarsening literature, different scoring functions have been defined: from reconstruction methods where the score represents the capability of reconstructing the original graph (see for example spatial coarsening [lefevre2010grass, riondato2017graph] and spectral coarsening [kumar2023featured]); to reconstruction-free method where the scoring function is measuring interesting features for the specific application (for a review see [hashemi2024comprehensive]). We define a scoring function that accounts for the balance between the topological characteristics distortion and the complexity of the coarsened graph.
@@
The scoring function is a combination of the proportion of remaining edges in the graph and the topological difference between the original and the spatial graph, measured via the bottleneck distance between persistent diagrams. To match the order of magnitude of both terms, $\lambda$ is set to the inverse of the maximal distance between the persistence diagrams of the initial and a reduced graph giving $\lambda(G)^{-1}:=\max_{\theta}d_{B}\left(PD(G),PD(f_{\theta}(G)))\right)$. Our topological spatial graph coarsening method minimizes the score with respect to $\theta$ to find a satisfying reduction level. It outputs the coarsened graph corresponding to this optimal $\theta$. In Figure 3, we show a spatial graph and its corresponding reduced graph obtained with our topological spatial graph coarsening (here $\theta^{*}=0.4$) using average node positioning. We also display the persistent diagrams of both graphs, showing that the persistent diagram of the reduced graph displays the main topological features (points far away from the diagonal) of the original one, while most of the “topological noise” (points close to the diagonal) have been removed, demonstrating the effect of the graph simplification.
@@ 5 Property of Spatial Graph Coarsening
We often want coarsening methods to be coherent and robust to certain modifications of the initial graph. In this work, we show that our topological spatial graph coarsening is equivariant with respect to certain group actions. In particular, as we are dealing with nodes embedded in euclidean space, the actions of interest are rotation, translation, reflection, and scaling. Let’s start by defining the actions on the spatial graph.
@@
Consider a spatial graph $G=(V,E,X)$ with coordinated in $X\in\mathbb{R}^{p\times n}$, $|V|=n$, and the similarity group $Sim(p)$. We can define a set of equivalent graphs $G]=\{\psi(G,(R,A,k))=(V,E,kRX+A),(R,A,k)\in Sim(p)\}$ by applying the group action via $\psi:\mathcal{G}\times Sim(p)\rightarrow\mathcal{G}$, where $R\in O^{p}(\mathbb{R})$ is an orthogonal matrix, $A\in\mathbb{R}^{p}$ is a vector translating the graph, and $k>0$ is a strictly positive constant value scaling the graph coordinates.
In the paper, we focus on graphs with nodes embedded in $p=2,3$ but the results holds in higher dimension. The symmetric group $Sym(p)$ is one of the common group studied in shape and image analysis. Its elements $(R,A,k)$ can rotate, scale, and translate the original object. In the case of spatial graphs, the equivalence class of the original graph $G$ contains all the graphs that share the same set of nodes $V$, the same connectivity structure $E$, and a similarity transformation of the spatial coordinated of the nodes in $\mathbb{R}^{p}$ (i.e. a set of rotated, scaled, and translated coordinates).
The objective of this section is to study the properties of the spatial graph coarsening procedure. While the writing might seems a bit cumbersome, the idea behind are simple and represented in an illustration in Figure 4. We want to show how the procedure is retrieving the same graph yet rotated, translated and scaled when an element of the symmetric group is applied to the original spatial graph. To do so, we proceed step by step. Firstly, we look at the spatial graph coarsening and we prove the parameter equivariance property (Section 5.1). Secondly, we look at the topological spatial graph coarsening. To prove its invariance under the group action, we need to look at how the action effect the scoring function (Section 5.2). Finally, we prove the main result of this section, Proposition 3, that shows the equivariance of our topological spatial coarsening procedure under the symmetric group action (Section 5.3).
@@ 5.1 Parameter equivariance property for the spatial graph coarsening
Consider a graph $G=(V,E,X)$ with node positions $X\in\mathbb{R}^{p},p=2,3$ and a length threshold $\theta\in\mathbb{R}_{+}$. The coarsening procedure introduced in Section 2 that defines the coarsened graph $f_{\theta}(G)=(f^{V}_{\theta}(G),f^{E}_{\theta}(G),f^{X}_{\theta}(G))$ only depends on the edge lengths $\ell_{u,v},(u,v)\in E$ of the original spatial graph $G$. It is trivial to see that the coarsening procedure is invariant under rotation and translation, as they are isometries and thus they leave the edge lengths unchanged. The coarsening procedure is also equivariant under scaling, if the parameter $\theta$ is properly rescaled. More precisely, let us consider $(k,R,A)\in Sim(p)$ where the scaling factor is $k\in\mathbb{R}_{+}$, the resulting graph $G^{\prime}=(V,E,kRX+A)$ has edge lengths $\ell^{\prime}_{u,v}=k\ell_{u,v}$. We want to prove that
@@
The coarsening function is fully equivariant under the Euclidean Group acting $\mathbb{E}(d)$, where there is no scaling component.
@@
Once we have proved the property of the coarsening procedure defined in Section 2, we would like to prove the equivariance property for the proposed topological spatial coarsening procedure of Section 4. For that, we first prove that the persistent diagram is invariant under rotation and translation of the object and it is equivariant under scaling.
@@
The invariance under rotation and translation has been proven in the literature of persistence diagrams [chazal2009proximity, chazal2014persistence]; the equivariance under scaling follows from the definition of persistence diagrams in the context of spatial graphs. More precisely, we will show that that each simplex appearing in the filtration of the original graph $G$ appears at a time scaled by $k$ in the filtration of the modified graph $\psi(G,(R,A,k))$. Therefore, each topological feature is born and died at scaled time, yielding the result. Therefore, we need to prove the result on simplices. In the following, let use consider the action of a similarity with only a scaling part, that is $G^{\prime}=\psi(G,(Id,0,k))$, where $k>0$. as we proved that the isometry part (rotation and translation) leaves the persistent diagram unchanged. Therefore, the scaled coordinates are given by $X^{\prime}=kX$ where $X$ are the Euclidean coordinates of the original graph $G$. Recall that the filtration $\mathcal{C}(G)=(\mathcal{C}_{r}(G))_{r\geq 0}$ is defined as the Vietoris-Rips filtration where the metric is given by the shortest-path distance over the graph (see Section 3.2). From the definition of this filtration, for each scale $r\geq 0$ we have that $\mathcal{C}_{kr}(G^{\prime})=\mathcal{C}_{r}(G)$, as all shortest-path distances in $G^{\prime}$ are the same as the ones in $G$ except that they are scaled by $k$. In the specific case of spatial graphs, in Section 3.2, we advocated the use of a modified filtration $\tilde{\mathcal{C}}$ to better capture cycles generated by triangles in the graph. To transpose the previous result to this modified filtration, we need to study the filtration arrival of the 2-simplices. Given a 2-simplex $(v_{1},v_{2},v_{3})$ that appears at scale $r$ in filtration $\tilde{\mathcal{C}}(G)$, let us check that it appears at scale $kr$ in filtration $\tilde{\mathcal{C}}(G^{\prime})$. Let $x_{1}$, $x_{2}$ and $x_{3}$ (resp. $x_{1}^{\prime}$, $x_{2}^{\prime}$ and $x_{3}^{\prime}$) represent the coordinates of $v_{1}$, $v_{2}$ and $v_{3}$ in $G$ (resp. $G^{\prime}$) (recall that this makes sense as $G$ and $G^{\prime}$ have the same set of nodes). From the definition of $\tilde{\mathcal{C}}(G^{\prime})$, the simplex $(v_{1},v_{2},v_{3})$ appears at scale:
@@ 5.3 Equivariance of the topological spatial graph coarsening
Given the two proposition above, we can now look at whole topological spatial graph coarsening procedure and show that is it equivariant under the symmetric group action.
@@ 6 Experiments and Real World Examples
We illustrate the proposed method on simulated and real world data. The simulated data consists of a spatial graph where the nodes are positioned on a ring with a random effect and the edges are sampled at random. We test the spatial graph coarsening procedure with average degree positioning. We then move to two real world dataset. We simplify the road network of Marseille, France using the degree positioning [boeing2025modeling]. We test the effect of the topological spatial graph coarsening procedure of a classification task for the dataset of network fungi [fricker2025fungi].
@@ 6.2 Single Spatial Graph Coarsening
We start by considering a single spatial graph generated in the following fashion. Consider $n=100$ points uniformly sampled on an annulus with outer radius $1$ and inner radius $0.7$. We include in the graph only a fraction $p=0.1$ of all the n(n1)/2n(n-1)/2 possible edges, those with smallest length (given by euclidean distance). In Figure 6 (left), we can see original and the reduced graph. The position of the nodes in the coarsened graph follow the average positioning procedure described in Section 2. The new nodes are positioned by computing the average of the merged nodes. The optimal reduction is chosen by looking at the score values in Figure 6 (right). We plot the score $S_{\theta}(G)=|E(\theta)|/|E|+\lambda\cdot d_{B}(PD(G),PD(G(\theta)))$ defined as in Equation (4.1) and its two components, the ratio of edges and the bottleneck distance scaled by $\lambda$ set as the inverse of the maximum of the bottleneck distance. The optimal threshold $\theta^{\star}$ is found at the minimum of the scoring function.
@@
As a real world example, we select the road network of Marseille, France (data publicly available in the OSMnx package [boeing2025modeling]). The graph is made of all the portions of road of the following type: motorway, motorway link, primary road, secondary road, secondary road link. Road networks are a perfect example of spatial graphs, with natural constrains on the road topology and with high complexity in terms of nodes and edges. The original and the reduced graphs are presented in Figure 8. The position of the nodes is computed using the degree positioning described in Section 2. For each new hypernode, we choose the position of the original node with the highest connectivity - i.e. higher degree. In this road network example, this is equivalent to assign the position of the crossing between the highest number of streets.

### v2

@@ 3.2 Persistence Diagram for Spatial Graphs
We want to adapt the above construction of persistence diagrams of point clouds data to spatial graphs data. To do so, we follow the same general idea: build a nested family of simplicial complexes on top of graphs and monitor how the topological components evolve through this filtration. Previous approaches for computing filtrations over graphs considered a family of nested subgraphs of $G$, $(G_{r})_{r\geq 0}$ (by filtering the edges based on weights or other graph characteristics like degree) and compute simplicial complexes over each subgraph $C(G_{r})$ where $C$ can be for example the clique complex (also called flag complex) or the neighborhood complex [lovasz1978kneser, aharoni2005eigenvalues, horak2009persistent]. These approaches differ from the point cloud pipeline on different aspects. First, the scale is given by a substructure of the initial object (here the graph) while for point clouds we consider the whole dataset and compute simplicial complexes at different scales. Secondly, some simplices are never included in the filtration (e.g., a 1-simplex that correspond to an edge that is not in $G$ will never be include in the clique complex).
We propose a new way of computing a filtration of simplicial complexes over a graph that is closer to the construction for point clouds. The essence of our approach lies in the fact that the VR filtration is entirely determined by the distances between all pairs of points in the data. Thus, we propose to replace the Euclidean distance used in point clouds by the shortest path distance over the graph, denoted by $d_{G}$ and defined for all pairs of vertices $u$ and $v$ by
@@
where we recall that the length $\ell_{u,v}$ of an edge $(u,v)$ is given by the Euclidean distance $d(x_{u},x_{v})$. Therefore, we define a simplicial complex at scale $r>0$ over the graph $G$ as $\mathcal{C}_{r}(G)=VR_{r}(V,d_{G})$.
@@
If the graph is the complete graph, the shortest path distance coincides with the Euclidean distance and we recover the classic $VR_{r}$ filtration.
The proposed filtration $\mathcal{C}(G)=(\mathcal{C}_{r}(G))_{r\geq 0}$ has the drawback that cycles formed by a triplet of edges (empty triangle) do not persist in the filtration. In fact, the 2-simplex (filled triangle) appears in the filtration exactly when the longest edge is included, making the cycle disappear at the same scale at which it appears. Thus, this type of cycles are not represented in the induced persistence diagram. As we would like to represent in the persistence diagram any cycles - no matter the number of edges - we propose to modify the construction of the simplicial complexes $\mathcal{C}_{r}(G)$ by enlarging the scale at which 2-simplices appear. This novel filtration is defined below and illustrated in Figure 2 in comparison with the classic Vietoris-Rips filtration.
@@
Figure 2: Illustration of the difference between the classic Vietoris-Rips filtration (top) and the proposed triangle-aware filtration (bottom), where the 2-simplex appears at scale $r=5$ and $r=\min\{3+4,\ 4+5,\ 3+5\}=7$, respectively.
@@ 4 Topological Spatial Graph Coarsening
Given a sequence of coarsened graphs $f_{\theta}(G)$ obtained by exploring the set $\Theta$, we define a scoring function to select the optimal coarsened graph. In graph coarsening literature, different scoring functions have been defined: from reconstruction methods where the score represents the capability of reconstructing the original graph (see for example spatial coarsening [lefevre2010grass, riondato2017graph] and spectral coarsening [kumar2023featured]); to reconstruction-free method where the scoring function is measuring interesting features for the specific application (for a review see [hashemi2024comprehensive]). We define a scoring function that accounts for the balance between the topological characteristics distortion and the complexity of the coarsened graph.
@@
The scoring function is a combination of the proportion of remaining edges in the graph and the topological difference between the original and the reduced graph, measured via the bottleneck distance between persistent diagrams. To match the order of magnitude of both terms, $\lambda$ is set to the inverse of the maximal distance between the persistence diagrams of the initial and a reduced graph giving $\lambda(G)^{-1}:=\max_{\theta}d_{B}\left(PD(G),PD(f_{\theta}(G)))\right)$. Our topological spatial graph coarsening method minimizes the score with respect to $\theta$ to find a satisfying reduction level. It outputs the coarsened graph corresponding to this optimal $\theta$. In Figure 4, we show a spatial graph and its corresponding reduced graph obtained with our topological spatial graph coarsening (here $\theta^{*}=0.4$) using average node positioning. We also display the persistent diagrams of both graphs, showing that the persistent diagram of the reduced graph displays the main topological features (points far away from the diagonal) of the original one, while most of the “topological noise” (points close to the diagonal) have been removed, demonstrating the effect of the graph simplification.
@@ 5 Property of Spatial Graph Coarsening
We often want coarsening methods to be coherent and robust to certain modifications of the initial graph. In this work, we show that our topological spatial graph coarsening is equivariant with respect to certain group actions. In particular, as we are dealing with nodes embedded in Euclidean space, the actions of interest are rotation, translation, reflection, and scaling. Let’s start by defining how these groups act on the spatial graph.
@@
Consider a spatial graph $G=(V,E,X)$ with coordinated in $X\in\mathbb{R}^{p\times n}$, $|V|=n$, and the similarity group $Sim(p)$. We can define a set of equivalent graphs $G]=\{\psi(G,(R,A,k))=(V,E,kRX+A),(R,A,k)\in Sim(p)\}$ by applying the group action via $\psi:\mathcal{G}\times Sim(p)\rightarrow\mathcal{G}$, where $R\in O^{p}(\mathbb{R})$ is an orthogonal matrix, $A\in\mathbb{R}^{p}$ is a vector translating the graph, and $k>0$ is a strictly positive constant value scaling the graph.
In the paper, we focus on graphs with nodes embedded in $\mathbb{R}^{p}$ with $p=2,3$ but the results hold in higher dimension. The symmetric group $Sym(p)$ is one of the common group studied in shape and image analysis. Its elements $(R,A,k)$ can rotate, scale, and translate the original object. In the case of spatial graphs, the equivalence class of the original graph $G$ contains all the graphs that share the same set of nodes $V$, the same connectivity structure $E$, and a similarity transformation of the spatial coordinated of the nodes in $\mathbb{R}^{p}$ (i.e. a set of rotated, scaled, and translated coordinates). The objective of this section is to study the properties of the spatial graph coarsening procedure. While the writing might seems a bit cumbersome, the idea behind are simple and represented in an illustration in Figure 5. We want to show how the procedure is retrieving the same graph yet rotated, translated and scaled when an element of the symmetric group is applied to the original spatial graph. To do so, we proceed step by step. Firstly, we look at the spatial graph coarsening and we prove the parameter equivariance property (Section 5.1). Secondly, we look at the topological spatial graph coarsening. To prove its invariance under the group action, we need to look at how the action effect the scoring function (Section 5.2). Finally, we prove the main result of this section, Proposition 3, that shows the equivariance of our topological spatial coarsening procedure under the symmetric group action (Section 5.3).
@@ 5.1 Parameter equivariance property for the spatial graph coarsening
Consider a graph $G=(V,E,X)$ with node positions $X\in\mathbb{R}^{p},p=2,3$ and a length threshold $\theta\in\mathbb{R}_{+}$. The coarsening procedure introduced in Section 2 that defines the coarsened graph $f_{\theta}(G)=(f^{V}_{\theta}(G),f^{E}_{\theta}(G),f^{X}_{\theta}(G))$ is governed by the edge lengths of the original spatial graph $G$. It is trivial to see that the coarsening procedure is invariant under rotation and translation, as they are isometries and thus they leave the edge lengths unchanged. The coarsening procedure is also equivariant under scaling, if the parameter $\theta$ is properly rescaled. More precisely, let us consider $(k,R,A)\in Sim(p)$ where the scaling factor is $k\in\mathbb{R}_{+}$, the resulting graph $G^{\prime}=(V,E,kRX+A)$ has edge lengths $\ell^{\prime}_{u,v}=k\ell_{u,v}$. We want to prove that
@@
The relationships between the elements in Proposition 1 are detailed in the diagram below:
@@
The coarsening function is fully equivariant under the Euclidean group actionn $\mathbb{E}(d)$, where there is no scaling component.
@@
Once we have proved the property of the coarsening procedure defined in Section 2, we would like to prove the equivariance property for the proposed topological spatial coarsening procedure defined in Section 4. For that, we first prove that the persistent diagram is invariant under rotation and translation of the object and it is equivariant under scaling.
@@
The invariance under rotation and translation has been proven in the literature of persistence diagrams [chazal2009proximity, chazal2014persistence]; the equivariance under scaling follows from the definition of persistence diagrams in the context of spatial graphs. More precisely, we want to show that that each simplex appearing in the filtration of the original graph $G$ at $r_{b}$, it appears at $kr_{b}$ in the filtration of the modified graph $\psi(G,(R,A,k))$. Therefore, each topological feature is born and died at scaled time, yielding the result. Therefore, we need to prove the result on simplices. In the following, let us consider the action of a similarity with only a scaling part, that is $G^{\prime}=\psi(G,(Id,0,k))$, where $k>0$, as we proved that the isometry part (rotation and translation) leaves the persistent diagram unchanged. Therefore, the scaled coordinates are given by $X^{\prime}=kX$ where $X$ are the Euclidean coordinates of the original graph $G$. Recall that the filtration $\mathcal{C}(G)=(\mathcal{C}_{r}(G))_{r\geq 0}$ is defined as the Vietoris-Rips filtration where the metric is given by the shortest-path distance over the graph (see Section 3.2). From the definition of this filtration, for each scale $r\geq 0$ we have that $\mathcal{C}_{kr}(G^{\prime})=\mathcal{C}_{r}(G)$, as all shortest-path distances in $G^{\prime}$ are the same as the ones in $G$ except that they are scaled by $k$. In the specific case of spatial graphs, in Section 3.2, we advocated the use of a modified filtration $\tilde{\mathcal{C}}$ to better capture cycles generated by triangles in the graph. To transpose the previous result to this modified filtration, we need to study the filtration arrival of the 2-simplices. Given a 2-simplex $(v_{1},v_{2},v_{3})$ that appears at scale $r$ in filtration $\tilde{\mathcal{C}}(G)$, let us check that it appears at scale $kr$ in filtration $\tilde{\mathcal{C}}(G^{\prime})$. Let $x_{1}$, $x_{2}$ and $x_{3}$ (resp. $x_{1}^{\prime}$, $x_{2}^{\prime}$ and $x_{3}^{\prime}$) represent the coordinates of $v_{1}$, $v_{2}$ and $v_{3}$ in $G$ (resp. $G^{\prime}$) (recall that this makes sense as $G$ and $G^{\prime}$ have the same set of nodes). From the definition of $\tilde{\mathcal{C}}(G^{\prime})$, the simplex $(v_{1},v_{2},v_{3})$ appears at scale:
@@ 5.3 Equivariance of the topological spatial graph coarsening
Given the two propositions above, we can now look at the whole topological spatial graph coarsening procedure and show that is it equivariant under the symmetric group action.
@@ 6 Experiments and Real World Examples
We illustrate the proposed method on simulated and real world data. The simulated data consists of a spatial graph where the nodes are positioned on a ring with a random effect and the edges are sampled at random. We test the spatial graph coarsening procedure with average degree positioning. We then move to two real world dataset. In the first, we simplify the road network of the city of Marseille (France) using the degree positioning [boeing2025modeling]. In the second, we test the effect of the topological spatial graph coarsening procedure of a classification task for a dataset of network fungi [fricker2025fungi].
@@ 6.2 Single Spatial Graph Coarsening
We start by considering a single spatial graph generated in the following fashion. Consider $n=100$ points uniformly sampled on an annulus with outer radius $1$ and inner radius $0.7$. We include in the graph only a fraction $p=0.1$ of all the n(n1)/2n(n-1)/2 possible edges, those with smallest length (given by Euclidean distance). In Figure 7 (left), we can see original and the reduced graph. The position of the nodes in the coarsened graph follow the average positioning procedure described in Section 2. The optimal reduction is chosen by looking at the score values in Figure 7 (right). We plot the score $S_{\theta}(G)=|E(\theta)|/|E|+\lambda\cdot d_{B}(PD(G),PD(G(\theta)))$ defined as in Equation (4.1) and its two components, the ratio of edges and the bottleneck distance scaled by $\lambda$ set as the inverse of the maximum of the bottleneck distance. The optimal threshold $\theta^{\star}$ is found at the minimum of the scoring function.
@@
As a real world example, we select the road network of the city of Marseille in south of France (the data are publicly available in the OSMnx package by [boeing2025modeling]). The graph is made of all the portions of road of the following type: motorway, motorway link, primary road, secondary road, secondary road link. Road networks are a perfect example of spatial graphs, with natural constrains on the road topology and with high complexity in terms of nodes and edges. The original and the reduced graphs are presented in Figure 9. The position of the nodes is computed using the degree positioning described in Section 2. For each new hypernode, we choose the position of the original node with the highest connectivity - i.e. higher degree. In this road network example, this is equivalent to assign the position of the crossing between the highest number of streets.